# Differentially Private Knowledge Distillation via Synthetic Text Generation

### v1

@@ Abstract
Large Language models (LLMs) are achieving state-of-the-art performance in many different downstream tasks. However, the increasing urgency of data privacy requires LLMs to train with Differential Privacy (DP) on private data. Concurrently it is also necessary to compress LLMs for real-life deployments on resource-constrained devices or latency-sensitive applications. Differential privacy and model compression generally must trade off utility loss to achieve their objectives. Moreover, concurrently achieving both can result in even more utility loss. To this end, we propose a novel differentially private knowledge distillation algorithm that exploits synthetic data generated by a differentially private LLM. The knowledge of a teacher model is transferred onto the student in two ways: one way from the synthetic data itself, the hard labels, and the other way by the output distribution of the teacher model evaluated on the synthetic data, the soft labels. Furthermore, if the teacher and student share a similar architectural structure, we can further distill knowledge by exploiting hidden representations. Our results show that our framework substantially improves the utility over existing baselines with strong privacy parameters, $\epsilon=2$, validating that we can successfully compress autoregressive LLMs while preserving the privacy of training data.
@@ 1 Introduction
Large Language Models (LLMs) have accomplished amazing feats, from generating human-like text to improving human productivity through AI-powered assistants. However, recent work has demonstrated that practical privacy attacks are possible due to LLMs’ memorization of training data Carlini et al. (2019, 2021). Differential Privacy (DP) is widely used to mitigate this type of information leakage Dwork (2006). To guarantee DP for machine learning models, DP is operationalized within the learning algorithm, known as DP-SGD Abadi et al. (2016). The standard training paradigm of LLMs is to first pre-train an LLM on large publicly available data, then fine-tune it on a specific downstream task using private data. Recent works in differentially private training have taken advantage of this paradigm by initializing an LLM using pre-trained weights and then applying DP-SGD in the fine-tuning process, which can significantly boost its performance Li et al. (2021); Yu et al. (2021); Ganesh et al. (2023).
@@
Due to the huge computational and memory demands that LLMs possess, it becomes necessary to include model compression in this training paradigm. For example, deployment of LLMs for client devices where memory resources are scarce, or real-time systems such as sentence completion in text editors where inference latency is crucial Xu et al. (2023). One popular technique for model compression that our work focuses on is task-specific Knowledge Distillation (KD) Hinton et al. (2015), which distills the knowledge of a larger teacher model onto a compressed student model. Although knowledge distillation in LLMs has been quite successful Jiao et al. (2019); Sanh et al. (2019); Sun et al. (2019), very little work has explored compressing LLMs with DP Mireshghallah et al. (2022); Yu et al. (2023a, b). Since achieving DP or KD for an LLM already trade-off utility loss, accomplishing both can result in substantial utility loss Mireshghallah et al. (2022), which highlights the challenge of this problem.
@@
Figure 1: A brief overview of our differentially private knowledge distillation framework which can be broken down in three steps: (1) A pre-trained teacher model is fine-tuned on a private corpus using DP-SGD. (2) Next, the teacher model will generate synthetic data using control codes. (3) Finally, a pre-trained student model is fine-tuned on the DP synthetic data with the DP fine-tuned teacher for knowledge distillation.
@@
The goal of our work is to enable efficient DP-compressed model deployment while minimizing model utility degradation. We introduce a simple, but effective, Differentially Private Knowledge Distillation framework, shown in Figure 1, to accomplish this. (1) We employ DP-SGD to fine-tune a teacher model on a training dataset, where the teacher accurately captures the distributional properties of the training data. (2) Then we generate synthetic text closely resembling the original data distribution without incurring additional privacy loss. (3) Lastly, the student model is trained using a non-private optimizer on the differentially private synthetic data while the teacher distills its knowledge to the student.
@@
In summary, our key contributions are the following: (1) We introduce a novel differentially private knowledge distillation algorithm for autoregressive LLMs using synthetic text generation. Our framework utilizes two forms of knowledge transfer from the teacher: one from the synthetic text data, and the other from the teacher output distribution. (2) We comprehensively demonstrate that our algorithm significantly improves the utility of the student model over existing baselines with strong privacy parameters, $\epsilon=2$; In particular, the alignment step between the output distribution of the teacher and student is a crucial component to boosting the performance of the student. (3) We perform extensive experiments exploring different hyperparameters to provide insight into our approach. Additionally, if the teacher and student share a similar architectural structure, we can exploit the hidden representation of the teacher and student to boost the performance of our framework even further.
@@ 2 Related Works
Mireshghallah et al. (2022) first introduced DP model compression and gave a framework for DP knowledge distillation. However, their framework requires using DP-SGD twice: once for the teacher and once for the student. This procedure leads to suboptimal results for text classification tasks. Wang et al. (2019) also explored private model compression, but is limited by the number of queries to the teacher for knowledge distillation. Yu et al. (2023b) explores DP model compression during the pre-training stage. Our work focuses solely on the private fine-tuning stage. Another line of work similar to our approach is Data-Free Distillation Chen et al. (2019); Haroush et al. (2020) where a generative model synthesizes data to substitute training data. Since no private data is used, this results in no privacy loss, $\epsilon=0$. Our approach uses the private dataset to generate synthetic data in a differentially private manner.
@@
DP Synthetic Data for Downstream Task. Promising progress has been made in generating high-fidelity, DP synthetic text Bommasani et al. (2019); Mattern et al. (2022); Kurakin et al. (2023); Yue et al. (2022). These works have substantially reduced the utility gap between DP-SGD on real data and SGD on DP synthetic data for classification tasks. One work investigated synthetic data generation from DP foundational models to train lightweight downstream models Yu et al. (2023a). However, our method differs from theirs by the following: (1) Their framework only considers knowledge transfer from the synthetic data. Our results demonstrate that other forms of knowledge distillation, in particular, the output distribution and hidden representations, are crucial to obtain strong performance for compressed models. (2) Their framework needs large foundational models to generate high-enough quality synthetic text data, which requires model parallelism on four V100 GPUs for fine-tuning even with LoRA Hu et al. (2021); Yu et al. (2021). And currently no open-source implementation exists for DP model parallelism. In contrast, our framework can obtain strong results without relying on large foundational models and thus does not need model parallelism.
@@ 3.1 Knowledge Distillation
Knowledge Distillation (KD) aims at transferring the knowledge of a teacher LLM $T$ to a smaller student LLM $S$. The standard approach is to train the student to mimic the output distribution of the teacher and the true labels. Let $P_{T}$ and $P_{S}$ be the softmax outputs of the teacher and student networks, respectively, $Z_{T}$ and $Z_{S}$ be the logit outputs of the teacher and student networks, respectively, and $\mathbf{y}$ be the true labels. The student is trained to minimize the linear combination of the supervised learning loss and the distillation loss:
$\mathcal{L}_{\text{KD}}=(1-\lambda)\mathcal{L}_{\text{CE}}(\mathbf{y},P_{S})+\\ \lambda t^{2}\mathcal{L}_{\text{KL}}(\sigma(Z_{T}/t),\log\left(\sigma(Z_{S}/t)\right))$ (1)
where $\lambda$ is a weight hyperparameter, $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, $\mathcal{L}_{\text{KL}}$ is the Kullback-Leibler divergence, $\sigma$ is the softmax function, and the second term is the distillation loss which follows the softmax-temperature using a temperature parameter $t$ Hinton et al. (2015). $t$ controls the smoothness of the output distribution, so higher values produce a softer probability distribution which allows the student to extract richer information. At inference time, $t$ is set to 1 to recover the standard softmax. y is known as the hard label because of its one-hot encoding while $\sigma(Z_{T}/t)$ is the soft label.
@@ 3.3 Differential Privacy
The standard approach to integrating DP for deep learning models is by modifying the stochastic gradient descent algorithm (SGD), called DP-SGD Abadi et al. (2016). The modifications involve calculating per-sample gradients, rather than per-batch gradients, clipping them if their l-2 norm exceeds some threshold, then adding noise sampled from a Gaussian distribution. In the following section, we will make use of DP-SGD.
@@ 4.1 Problem Formulation
We now describe our framework in detail, which is summarized at a high level in Algorithm 1. Given a task-specific dataset $D$, a pre-trained, generative teacher model $T$, a pre-trained, generative student model $S$, and a privacy budget $(\epsilon,\delta)$, our goal is to produce a student model such that $S$ minimizes Equation 1, the size of $S$ is significantly smaller than $T$, and $S$ is $(\epsilon,\delta)$-DP with respect to $D$. We will also assume access to a set of control codes $C$ Keskar et al. (2019) in our solution.
Parameter DP-SGD DPKD DP Syn Ours
Teacher
Epochs - 101, 202 101, 202 101, 202
Learning Rate - 1e-41, 2 1e-41, 2 1e-41, 2
Batch Size - 40961, 2 40961, 2 40961, 2
Clipping Norm - 1.01, 2 1.01, 2 1.01, 2
Student
Epochs 251, 402 251, 402 51, 32 121, 2
Learning Rate 1e-41, 2 1e-41, 2 1e-61, 2e-52 8e-51, 2
Batch Size 40961, 2 40961, 2 161, 2 161, 2
Clipping Norm 1.01, 2 1.01, 2 - -
Table 1: Training hyperparameters for the teacher and student model of every method for the 1Yelp and 2Big Patent datasets.
The DPKD framework introduced in Mireshghallah et al. (2022) tackles this problem setup by first fine-tuning a teacher model with DP-SGD, and then using DP-SGD a second time to minimize Eq 1 for the student. This solution has three limitations: (1) using DP-SGD on the student model results in additional error proportional to the total number of parameters of the student model, in the worst case Li et al. (2022); (2) allocating the privacy budget between the teacher and student can make it difficult for the student to obtain useful signal for strict privacy settings, smaller $\epsilon$ values; (3) larger runtime and memory consumption with an additional use of DP-SGD.
Rather than applying DP-SGD twice, we take the view that only applying DP-SGD on the teacher model leads to a more optimal solution. The framework from Yu et al. (2023a) follows this approach, but, as our results will show, boosting the performance of the student requires the inclusion of more information to distill knowledge from the teacher to the student.
Algorithm 1 Differentially Private Knowledge Distillation via Synthetic Data.
Input: A pre-trained teacher model $T$, pre-trained student model $S$, private data $D$, control codes $C$, privacy budget $(\epsilon,\delta)$
@@
1:Train $T$ on $D$ using DP-SGD with privacy budget $(\epsilon,\delta)$-DP
2:Generate synthetic data $\tilde{D}$ using $T$ and $C$
3:Train $S$ on $\tilde{D}$ to minimize Equation 1
4:return $S$
4.2 Knowledge Distillation with DP Synthetic Text Data
@@
In our approach, we fine-tune a generative LLM using DP-SGD as our teacher model $T$, i.e., $T$ is $(\epsilon,\delta)$-DP. Afterward, $T$ is used to generate synthetic text data. In practice, we can use any synthetic text generation methodology, but we choose to follow an approach from Yue et al. (2022). In particular, we will utilize control codes Keskar et al. (2019) to provide more explicit control over text generation. Hence, sampling the next token $\hat{x}_{t+1}$ involves conditioning the model with a control code $c$ from a set of control codes $C$, in addition with the context vector $\mathbf{x}_{t}$:
$\hat{x}_{t+1}\sim p(x_{t+1}|\mathbf{x}_{t},c).$
@@
Control codes provide useful guidance to the model when we want the model to focus on specific topics or features when generating text. For example, when generating reviews, we can prompt the model with additional information about the product type and review score. A potential control code could look like the following: "Product type: Toy; Review score: 3.0." Models learn to use these control codes in their text generations when the codes are included in the training dataset. Using our working example, the model should be equipped to generate reviews for a Toy with a review score of 3.0 after fine-tuning. For our use case, we prepend each training sample with a control code containing categorical information about the sample. Then after training, we generate text data using only the control codes as the initial context vector from the training dataset, preserving the control code distribution.
@@
Once the synthetic data $\tilde{D}$ has been generated, we fine-tune the pre-trained student model $S$ on $\tilde{D}$ along with knowledge distilled from the teacher model $T$ to minimize Equation 1 without DP-SGD. With this setup, we can capture the knowledge of the teacher, which is a function of the entire private dataset, in two ways: the synthetically generated data and the output distribution of the teacher. Synthetic data provides a coarse way to transfer the knowledge of a teacher model, since one-hot encodings only give one piece of information about the output of the teacher. However, the knowledge distilled from the entire output distribution of the teacher is richer, providing additional information such as prediction uncertainty of the teacher that the student can learn from. Since the synthetic data closely resembles the original private data, minimizing Equation 1 using $\tilde{D}$ gives an accurate approximation to Equation 1 using $D$.
@@
Another advantage of this framework is that the synthetic text generator and teacher model are the same. Hence, we do not need to split the privacy budget and instead can give it entirely to the teacher model. This allows for better performance for the teacher, resulting in higher-quality text generation and richer information from the output distribution. Furthermore, since our method only requires DP-SGD once, which is for the teacher, we further reduce additional memory and computational costs compared to DPKD.
@@
Lastly, we designed our framework to be model agnostic, making no architectural assumptions between the teacher and student. However, in section 5.4, we show that our framework can also be improved by loosening this assumption. More specifically, we can add loss terms to Eq. 1 that involve aligning the hidden representations of the teacher and student model. Our method is precisely summarized in Algorithm 1.

### v2

@@ Abstract
Large Language models (LLMs) are achieving state-of-the-art performance in many different downstream tasks. However, the increasing urgency of data privacy puts pressure on practitioners to train LLMs with Differential Privacy (DP) on private data. Concurrently, the exponential growth in parameter size of LLMs necessitates model compression before deployment of LLMs on resource-constrained devices or latency-sensitive applications. Differential privacy and model compression generally must trade off utility loss to achieve their objectives. Moreover, simultaneously applying both schemes can compound the utility degradation. To this end, we propose DistilDP: a novel differentially private knowledge distillation algorithm that exploits synthetic data generated by a differentially private teacher LLM. The knowledge of a teacher LLM is transferred onto the student in two ways: one way from the synthetic data itself– the hard labels, and the other way by the output distribution of the teacher evaluated on the synthetic data– the soft labels. Furthermore, if the teacher and student share a similar architectural structure, we can further distill knowledge by aligning the hidden representations between both. Our experimental results demonstrate that DistilDP can substantially improve the utility over existing baselines, at least $9.0$ PPL on the Big Patent dataset, with strong privacy parameters, $\epsilon=2$. These promising results progress privacy-preserving compression of autoregressive LLMs. Our code can be accessed here: https://github.com/james-flemings/dp_compress.
@@ 1 Introduction
Large Language Models (LLMs) have accomplished amazing feats, from generating human-like text to improving human productivity through AI-powered assistants. However, recent work has demonstrated that practical privacy attacks are possible due to memorization of training data from LLMs Carlini et al. (2019, 2021). Differential Privacy (DP), a rigorous mathematical framework, is widely used to mitigate this type of information leakage Dwork (2006). To guarantee DP for machine learning models, DP is operationalized within the learning algorithm, known as DP-SGD Abadi et al. (2016). The standard training paradigm of LLMs is to first pre-train an LLM on large publicly available data, then fine-tune it on a specific downstream task using private data. Recent works in differentially private training have taken advantage of this paradigm by initializing an LLM using pre-trained weights and then applying DP-SGD in the fine-tuning process, which can significantly boost its performance Li et al. (2021); Yu et al. (2021); Ganesh et al. (2023), albeit with some reduction in model utility.
@@
Due to the huge computational and memory demands of LLMs, it becomes necessary to consider model compression in this training paradigm. For example, model compression is required when deploying LLMs on client devices where memory resources are scarce, or real-time systems such as sentence completion in text editors where inference latency is crucial Xu et al. (2023). One popular technique for model compression that our work focuses on is task-specific Knowledge Distillation (KD) Hinton et al. (2015), which distills the knowledge of a larger teacher model onto a compressed student model. Although knowledge distillation in LLMs has been quite successful Jiao et al. (2019); Sanh et al. (2019); Sun et al. (2019), few works have explored compressing LLMs with DP Mireshghallah et al. (2022); Yu et al. (2023a, b). Since achieving DP or KD for an LLM already trade-offs utility loss, accomplishing both can result in substantial utility loss, which highlights the challenge of this problem.
@@
Figure 1: A brief overview of our differentially private knowledge distillation framework, DistilDP, which can be broken down in three steps: (1) A pre-trained teacher model is fine-tuned on a private corpus using DP-SGD. (2) Next, the teacher model will generate synthetic data using control codes. (3) Finally, a pre-trained student model is fine-tuned on the DP synthetic data with the DP fine-tuned teacher for knowledge distillation.
A naive solution to this problem would be to apply DP-SGD when training the teacher model, then apply DP-SGD again during the knowledge distillation between the teacher and student. Since the teacher model is a function of the entire private corpus, applying DP-SGD during the fine-tuning of the teacher is seemingly unavoidable. Despite the additive computational and memory costs due to an additional run of DP-SGD, this solution has also been shown to have worst utility compared to differentially private fine-tuning the student without knowledge distillation Mireshghallah et al. (2022).
@@
The goal of our work is to enable efficient differentially private compression of LLM deployment while minimizing model utility degradation. To this end, we introduce DistilDP: a simple, but effective, Differentially Private Knowledge Distillation framework, shown in Figure 1. The innovation of our approach is that it avoids running an additional DP-SGD during the knowledge distillation by utilizing DP synthetic data. Our frameworks proceeds as follows: (1) First, DP-SGD is employed to fine-tune a teacher model on a training dataset, where the teacher accurately captures the distributional properties of the training data. (2) Then we generate synthetic text closely resembling the original data distribution without incurring additional privacy loss. (3) Lastly, the student model is trained using a non-private optimizer on the differentially private synthetic data while the teacher distills its knowledge to the student.
@@
In summary, our key contributions are the following: (1) We introduce DistilDP, a novel differentially private knowledge distillation algorithm for autoregressive LLMs using synthetic text generation. Our framework utilizes two forms of knowledge transfer from the teacher: one from the synthetic text data, and the other from aligning the teacher output distribution with the student. (2) We comprehensively demonstrate that our algorithm significantly improves the utility of the student model over existing baselines with strong privacy parameters, $\epsilon=2$; In particular, the alignment step between the output distribution of the teacher and student is a crucial component to boosting the performance of the student. (3) We perform extensive experiments exploring different hyperparameters to provide insight into our approach. Additionally, if the teacher and student share a similar architectural structure, we can align the hidden representation of the teacher and student to further boost the utility of our framework.
@@ 2 Related Works
DP-SGD Abadi et al. (2016) is the most widely used technique to guarantee DP for LLMs. Specifically, a pre-trained LLM trained on large amounts of public text data scraped from the internet is further fine-tuned on a private corpus $D$ using DP-SGD McMahan et al. (2017); Yu et al. (2021); Li et al. (2021). An orthogonal approach investigated differentially private decoding of LLMs Flemings et al. (2024); Majmudar et al. (2022); Ginart et al. (2022), thereby avoiding the use of DP-SGD.
Mireshghallah et al. (2022) first introduced DP model compression and gave a framework for DP knowledge distillation. However, their framework requires using DP-SGD twice: once for the teacher and once for the student. This procedure leads to suboptimal results for text classification tasks. Wang et al. (2019) also explored private model compression, but is limited by the number of queries to the teacher for knowledge distillation. Yu et al. (2023b) explores DP model compression during the pre-training stage; our work focuses solely on compression during fine-tuning. Another line of work similar to our approach is Data-Free Distillation Chen et al. (2019); Haroush et al. (2020) where a generative model synthesizes data to substitute training data. Since no private data is used, this results in no privacy loss, $\epsilon=0$. Our approach uses a private corpus to generate synthetic data in a differentially private manner.
@@
DP Synthetic Data for Downstream Task. Promising progress has been made in generating high-fidelity, DP synthetic text Bommasani et al. (2019); Mattern et al. (2022); Kurakin et al. (2023); Yue et al. (2022). These works have substantially reduced the utility gap between DP-SGD on real data and SGD on DP synthetic data for classification tasks.
Yu et al. (2023a) investigated synthetic data generation from DP foundational models to train lightweight downstream models. However, our method differs from theirs by the following: (1) Their framework only considers knowledge transfer from the synthetic data. Our results demonstrate that other forms of knowledge distillation, in particular, the output distribution and hidden representations, are crucial to obtain strong performance for compressed models. (2) Their framework needs large foundational models to generate high-enough quality synthetic text data, which requires model parallelism on four V100 GPUs for fine-tuning even with LoRA Hu et al. (2021); Yu et al. (2021). And to the best of our knowledge, there is no open-source implementation of DP model parallel training. In contrast, our framework can obtain strong results without relying on large foundational models and thus does not need model parallelism.
@@ 3.1 Knowledge Distillation
Knowledge Distillation (KD) aims at transferring the knowledge of a teacher LLM $T$ to a smaller student LLM $S$. The standard approach is to train the student to mimic the output distribution of the teacher and the true labels. Let $f_{T}(x)$ and $f_{S}(x)$ be the logit outputs of the teacher and student networks evaluated on an input data $x$, respectively. $P_{T}$ and $P_{S}$ be the softmax outputs of the teacher and student networks, respectively, $Z_{T}$ and $Z_{S}$ be the logit outputs of the teacher and student networks, respectively, and $\mathbf{y}$ be the true labels. The student is trained to minimize the linear combination of the supervised learning loss and the distillation loss:
$\mathcal{L}_{\text{KD}}(T,S,D)=\sum_{(x,\textbf{y})\in D}(1-\lambda)\mathcal{L}_{\text{CE}}(\mathbf{y},\sigma(f_{S}(x)))$
$+\lambda t^{2}\mathcal{L}_{\text{KL}}(\log(\sigma(f_{T}(x)/t)),\log\left(\sigma(f_{S}(x)/t)\right))$ (1)
where $\lambda$ is a weight hyperparameter, $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, $\mathcal{L}_{\text{KL}}$ is the Kullback-Leibler divergence, $\sigma$ is the softmax function, and the second term is the distillation loss which follows the softmax-temperature using a temperature parameter $t$ Hinton et al. (2015). At inference time, $t$ is set to 1 to recover the standard softmax. y is known as the hard label because of its one-hot encoding while $\sigma(f_{T}(x)/t)$ is the soft label.
@@ 3.3 Differential Privacy
The standard approach to integrating DP for deep learning models is by modifying the stochastic gradient descent algorithm (SGD), called DP-SGD Abadi et al. (2016). The modifications involve calculating per-sample gradients, rather than per-batch gradients, clipping them if their l-2 norm exceeds some threshold, then adding noise sampled from a Gaussian distribution. Note that per-sample gradient calculations typically reduce the parallel efficiency of GPUs; hence, applying DP-SGD has a negative impact on the training speed Yousefpour et al. (2021). In the following section, we will make use of DP-SGD.
@@ 4.1 Problem Formulation
Given a task-specific private dataset $D$, a pre-trained, generative teacher model $T$, a pre-trained, generative student model $S$, and a privacy budget $(\epsilon,\delta)$, our goal is to produce a student model such that $S$ minimizes Equation 3.1, the size of $S$ is significantly smaller than $T$, and $S$ is $(\epsilon,\delta)$-DP with respect to $D$. We will also assume access to a set of control codes $C$ Keskar et al. (2019) in our solution.
The DPKD framework introduced in Mireshghallah et al. (2022) tackles this problem setup by first fine-tuning a teacher model with DP-SGD, and then using DP-SGD a second time to minimize Eq 3.1 for the student. This solution has three limitations: (1) using DP-SGD on the student model results in additional error proportional to the total number of parameters of the student model, in the worst case Li et al. (2022); (2) allocating the privacy budget between the teacher and student can make it difficult for the student to obtain useful signal for strict privacy settings, smaller $\epsilon$ values; (3) larger runtime and memory consumption with an additional use of DP-SGD.
Rather than applying DP-SGD twice, we take the view that only applying DP-SGD on the teacher model leads to a more optimal solution. The framework from Yu et al. (2023a) follows this approach, but, as our results will show, the inclusion of more information about the teacher, such as the output distribution, substantially improves the utility of the student.
Algorithm 1 DistilDP: Differentially Private Knowledge Distillation via Synthetic Data
Input: A pre-trained teacher model $T$, pre-trained student model $S$, private data $D$, control codes $C$, max length of synthetic examples $T_{\text{max}}$, total number of synthetic examples $N$, privacy budget $(\epsilon,\delta)$, mixing parameter $\lambda$, temperature parameter $t$
@@
1:Preprocessing Step: Prepend control codes $C$ onto the dataset $D$, i.e., $D^{\prime}[i]=C[i]+D[i]$ $\forall i\in[|D|]$.
@@
2:Fine-tune the teacher $T$ on the preprocessed data $D^{\prime}$ using DP-SGD with privacy budget $(\epsilon,\delta)$-DP
3:Subsample a subset of the control codes $\tilde{C}\subseteq C$ where $|\tilde{C}|=N$
4:for $c\in\tilde{C}$ do
5:     Synthetic example $\tilde{d}\sim\prod\limits_{t=1}^{T_{\text{max}}}T(x_{t}|x_{1:t},c)$
6:     $\tilde{D}=\tilde{D}\cup\tilde{d}$
7:end for
8:Train Student $S=\min\limits_{S^{\prime}}\mathcal{L}_{\text{KD}}(T,S^{\prime},\tilde{D})$
9:return $S$
4.2 DistilDP: Knowledge Distillation with DP Synthetic Text Data
@@
We now introduce DistilDP, described in Algorithm 1. DistilDP fine-tunes a pre-trained, generative LLM using DP-SGD as our teacher model $T$, i.e., $T$ is $(\epsilon,\delta)$-DP. Afterward, $T$ is used to generate synthetic text data. In practice, we can use any synthetic text generation methodology, but we choose to follow an approach from Yue et al. (2022). In particular, we will utilize control codes Keskar et al. (2019) to provide more explicit control over text generation. Hence, sampling the next token $\hat{x}_{t+1}$ involves conditioning the model with a control code $c$ from a set of control codes $C$, in addition with the context vector $\mathbf{x}_{t}$: $\hat{x}_{t+1}\sim p(x_{t+1}|\mathbf{x}_{t},c)$.
@@
Control codes guide the model to focus on specific topics or features during text generation. For example, when generating reviews, we can prompt the model with additional information about the product type and review score. A potential control code could look like the following: "Product type: Toy; Review score: 3.0." Models learn to use these control codes in their text generations when the codes are included in the training dataset. Using our working example, the model should be equipped to generate reviews for a Toy with a review score of 3.0 after fine-tuning. For our use case, we prepend each training sample with a control code containing categorical information about the sample. Then after training, we generate text data using only the control codes as the initial context vector from the training dataset, preserving the control code distribution.
@@
Once the synthetic data $\tilde{D}$ has been generated, we fine-tune the pre-trained student model $S$ on $\tilde{D}$ along with knowledge distilled from the teacher model $T$ to minimize Equation 3.1 without DP-SGD. With this setup, we can capture the knowledge of the teacher, which is a function of the entire private dataset, in two ways: the synthetically generated data and the output distribution of the teacher. Synthetic data provides a coarse way to transfer the knowledge of a teacher model, since one-hot encodings only give one piece of information about the output of the teacher. However, the knowledge distilled from the entire output distribution of the teacher is richer, providing additional information such as prediction uncertainty of the teacher that the student can learn from. Since the synthetic data closely resembles the original private data, minimizing Equation 3.1 using $\tilde{D}$ gives an accurate approximation to Equation 3.1 using $D$.
@@
Another advantage of DistilDP is that the synthetic text generator and teacher model are the same. Hence, we do not need to split the privacy budget and instead can allocate it entirely to privately fine-tuning the teacher model. This allows for better performance for the teacher, resulting in higher-quality text generation and richer information from the output distribution. Furthermore, since our method only requires DP-SGD once, which is for the teacher, we further reduce additional memory and computational costs compared to DPKD.
@@
Lastly, we designed our framework to be model agnostic, making no architectural assumptions between the teacher and student. However, in section 5.4, we show that our framework can also be improved by loosening this assumption. More specifically, we can add loss terms to Eq. 3.1 that involve aligning the hidden representations of the teacher and student model. Our method is precisely summarized in Algorithm 1.